the decomposition of so2cl2 with a rate

CHM 1046 2. The decomposition ofsoacl. s first order in sonch with a rate constant of 1.42x10 at a given temperature. b. lf the initial concentration of sozcizis0.150 M, what is the concentration after 300 seconds? 3. The decomposition of substance A is found to be second order at 700 C with a rate constant of 7.02x103M st If the initial concentration of A is 0.125 M, what is the concentration of A after 350 seconds?
Need help. Don't understand
Answer #1

2a) Integrated form of first orde rate equation is

ln[A]t = -kt + ln[A]0

Where, k = rate constant , 1.42 × 10^-4/s

ln[A]t = concentration at time t , 0.65M

ln[A]0 = concentration at initial, 1M

t = time , ?

Substituting the values

ln(0.65M) = - (1.42 ×10^-4/s)×t + ln(1M)

t = - ln(0.65) s /1.42 × 10^-4

= 3034 seconds

2b) Initial concentration , [A]0 is 0.150MM

[A]t = ? , t = 30seconds , k=1.42×10^-4/s

Substituting the values,

ln[A] = - (1.42 × 10^-4/s) × 300 s + ln(0.150M)

2.303log[A] = - 0.0426 - 1.8975M

[A] = 0.144M

3.Integrated form of second order equation is

1/[A]t = kt + 1/[A]0

k =7.02 × 10^-3 /M s

[A]0 = 0.125M , t = 350 seconds

[A]t = ?

Substituting the values

1/[A]t = ( 7.02 ×10^-3/M s )× 350s + 1/0.125M

= 0.0956M